OpenStudy (anonymous):

Using the formula P = 2(a + b), find a when P = 15 and b = 5. 2.5 5 8 10

OpenStudy (anonymous):

so, fill in the numbers for the letters. 15=2(a+5) then 15=2a+10. subtract 10 from both sides...so 15-10=2a or 5=2a. then divide by 2. So 5/2=a. Which is 2.5